A firm manufactures two products; the net profit on product 1 is Birr 3 per unit and Birr 5 per unit on product 2. The manufacturing process is such that each product has to be processed in two departments D1 and D2. Each unit of product1 requires processing for 1 minute at D1 and 3 minutes at D2; each unit of product 2 requires processing for 2 minutes at D1 and 2 minutes at D2. Machine time available per day is 860 minutes at D1 and 1200 minutes at D2. How much of product 1 and 2 should be produced every day so that total profit is maximum. (solve with graphical method) Q3. Solve the following LP problem by graphical method: Maximise Z = 300X1 + 700X2 Subject to the constraints: X1 + 4X2 ≤ 20 2X1 + X2 ≤ 30 X1 + X2 ≤ 8 And X1, X2 ≥ 0 Q4.

1
Expert's answer
2022-02-03T17:31:38-0500

Solution:

Let "x_1" and "x_2" be levels of production of two products, then

"Z=3 x_1+5 x_2" is profit function and it should be maximized.

Subject to the constraints:

"x_1+2 x_2\le860;\\
3 x_1+2 x_2\le1200;\\"

"x_1\ge0,x_2\ge0"


Corners points are (0,430), (170,345), (400,0).

At (0,430), "Z=3(0)+5 (430)=2150"

At (170,345), "Z=3(170)+5 (345)=2235"(Maximum)

At (400,0), "Z=3(400)+5 (0)=1200"

Hence, the optimal solution to the given LP problem is : x1=170,x2=345 and max Z=2235.

2. Maximise "Z = 300x_1 + 700x_2" Subject to the constraints:

"x_1 + 4x_2 ≤ 20 
\\2x_1 + x_2 ≤ 30 
\\x_1 + x_2 ≤ 8 
\\x_1, x_2 ≥ 0"



Corner points are (0,5), (4,4), (8,0).

At (0,5), "Z = 300(0) + 700(5)=3500"

At (4,4), "Z = 300(4) + 700(4)=4000"(maximum)

At (8,0), "Z = 300(8) + 700(0)=2400"

Thus, at "x_1=4,x_2=4, Z=4000" is the maximum value.